OpenStudy (anonymous):

Sam wants to buy 8 pencils and 6 pens, which cost a total of $8.50. He realizes he doesn't have enough cash and instead takes only 4 of each, which costs only $5. What is the price of each pen and pencil?

OpenStudy (anonymous):

Pens are 75 cents, and pencils are 50 cents

OpenStudy (anonymous):

8p+6i=8.5 4p+4i=5 4p=5-4i p=1.25-i 8(1.25-i)+6i=8.5 10-8i+6i=8.5 10-2i=8.5 -2i=-1.5 i=.75 4p+4(.75)=5 4p+3=5 4p=2 p=.5

OpenStudy (anonymous):

Let the price of one pencil and one pen be x and y dollars respectively. Therefore from first condition, we have: 8x +6 y = 8.5--------(1) Now from second condition, we have: 4x + 4y = 5--------(2) Multiplying eq 2 by 2 and then subtracting it from (1), we get. 8x +6 y -(8x + 8y) = 8.5- 10 8x +6 y -8x - 8y = -1.5 -2y = -1.5 y= -1.5/-2= 1.5/2= 0.75 from eq (2) we get 4x + 4*0.75 = 5 4x + 3 = 5 4x = 5-3 4x = 2 x= 2/4= 0.5 Thus the price of each pencil and pen is $0.5 and $0.75 respectively.